OpenStudy (anonymous):

Why is the answer all real numbers? 9(n+3) = 3(9 + 3n) subtract 3n from both sides 6n + 27 = 27 subtract 27 from both sides 6n = 0 divide by 6 n = 0/6 n=0 So how could the answer be "all real numbers"? Thank you.

hartnn (hartnn):

on right side 3(9+3n) will give 27 + 9n too!

hartnn (hartnn):

3(9+3n) = 3*9 + 3*3n = 27 +9n

hartnn (hartnn):

got your error ?

hartnn (hartnn):

@GaiginJen

OpenStudy (yamyam70):

@hartnn GOKKKKKKKKKUUUUUUUUUUUUUU

OpenStudy (anonymous):

I am not asking the question properly. I will try to understand what I am doing wrong thank you for trying to help.

OpenStudy (yamyam70):

and by the way please do read the code of conduct found here and welcome to open study! http://openstudy.com/code-of-conduct :)

OpenStudy (anonymous):

Please, how did I violate the code of conduct? I thanked the person. I think I am in the wrong place.

OpenStudy (yamyam70):

Man I'm just welcoming you here, no violations :D